OpenStudy (anonymous):

Evaluate the integral: Sqrt[4-x^2], n= 0 to 2

OpenStudy (anonymous):

That's beatiful :-)

OpenStudy (anonymous):

use a trig. substitution. theta=2sintheta

OpenStudy (anonymous):

@bmp Lmao, what?

OpenStudy (anonymous):

The answer to it... Anyway, use x = 2sin(u) and dx = 2cos(u)du. The square root becomes: \[ \sqrt{4-x^2} = \sqrt{4-4\sin^2{u}} = 2\cos(u) \]

OpenStudy (blockcolder):

Not necessary. Think of it as an area of a quarter circle with radius 2.

OpenStudy (anonymous):

Using polar coordinates becomes too easy :-). But, yes, I agree with @blockcolder . That's a lot easier.

OpenStudy (anonymous):

\[\sqrt{a ^{2}+x ^{2}} use x=asine(\theta)\] \[\sqrt{a ^{2}-x ^{2}} use x=atan(\theta)\] \[\sqrt{x ^{2}-a ^{2}} use x=asec(\theta)\]

OpenStudy (anonymous):

Doesn't make the answer less pretty, tho. :-)

OpenStudy (anonymous):

im sure the teacher want to see trig subs, this is for his finals lol =) just saying

OpenStudy (anonymous):

Well then: Hint: Write \[ \int 4\cos^2{(u)}du = 4 \int (\frac{1}{2} \cos(2u) + \frac{1}{2})du\]and integrate term by term.

OpenStudy (anonymous):

TYPO---- last x=asec(theta),, its + not -..... !!!!!

OpenStudy (anonymous):

How on earth do u find this stuff fun?!!? @bmp I'll stick to what you wrtoe before.

OpenStudy (anonymous):

omg my signs is asll wrong lol ignroe it. ill rewrite it

OpenStudy (anonymous):

You dont have to. I got it.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Calculus is, very likely, the most fun subject as an undergrad for me.

OpenStudy (anonymous):

Lol, genius. I just want to finish this course and change my major

OpenStudy (anonymous):

No, not at all. I suck at Linear Algebra, for instance. I barely passed it. Change it to Mathematics :-) Just kidding.

OpenStudy (anonymous):

lmao! Yeaaah, that was definitely good joke.

OpenStudy (anonymous):

even my grammar is messed up now

OpenStudy (anonymous):

My native language is messed up too. I think it's a consequence of doing something related to an exact sciences.